5 yrs ago the age of peter was 4 times the age of his son. If the sum of the present age is 60 yrs find the peters present age.
Answer by ewatrrr(24785)

x son's age, (60-x) peter's age NOW
5yrs ago...
(60-x) - 5 = 4(x-5)
55-x = 4x - 20
75 = 5x
x = 15, son's age. Peter 45.
checking
5yrs ago...
40 = 4*10 CHECKS!